2. Suppose that X1 and X2 constitute a sample of size 2 from a popula- tion in which a typical value X is equal to either 1 or 2 with respective
probabilities
P{X=1}=0.7 P{X=2}=0.3
(a) ComputeE[X].
(b) ComputeVar(X).
(c) WhatarethepossiblevaluesofX=(X1+X2)/2?
(d) DeterminetheprobabilitiesthatXassumesthevaluesin(c).
(e) Using(d),directlycomputeE[X]andVar(X).
(f) Are your answers to (a), (b), and (e) consistent with the formulas
presented in this section?
Answer #1

(a) E(X) = E(X1) + E(X2) = 0.7 + 0.3 = 1

(b) Var (X) = (1-0.7)2 + (1-0.3)2 = 0.58

(c) possible values = (1+2)/2 = 1.5

(d) The probabilities are 0.7 for X=1 and 0.3 for X=2

(e) E(X) = 0.7+ 0.3 = 1

Var (X) = (1-0.7)2 + (1- 0.3 )2 = 0.58

(f) Yes, the answers in (a), (b) and (e) consistent with the formulas present in this section.